The supplement of an angle is 36 degrees less than twice the supplement of the complement of the angle. Find the measure of the supplement. Can you also explain it too? Thank you so much
Found 3 solutions by haileytucki, mananth, stanbon:
Answer by haileytucki(390)

Ok, let me know if this helps;-)
1. Let x=measure of the angle.
Two angles angles are supplementary is the sum of their measures is 180 degrees.
Two angles angles are complementary is the sum of their measures is 90 degrees.
supplement of the angle
= 180-x
complement of the angle
= 90-x
supplement of the complement of the angle
= 180-(90-x)=90+x
twice the supplement of the complement of the angle
= 2(90+x)=180+2x
The supplement of an angle is 36 degrees less than twice the supplement of the complement of the angle
=> (180-x) + 36 = 180+2x
Solve for x. You should get x=12.
Since we are looking for the supplement, then the answer is 180-x=180-12=168.

Answer by mananth(16949)

let the angle be x degrees
the supplement of this angle will be 180-x
..
The complement of the angle will be 90-x
supplement of this complement angle will be 180 - (90-x)
Twice this = 2[180-(90-x)]
Less 36
=2[180-(90-x)]-36
..
180-x = 2[180-(90-x)]-36
...
180-x= 2[180-90+x]-36
180-x=2[90+x]-36
180-x=180+2x-36
3x=180-180+36
3x=36
x= 12
supplement of 12 = 180-12 = 168

Answer by stanbon(75887)

The supplement of an angle is 36 degrees less than twice the supplement of the complement of the angle. Find the measure of the supplement. Can you also explain it too?
-----------------------
angle = x
supplement = 180-x
complement = 90-x
supplement of the complement = 180-(90-x) = 90+x
---------------------
Equation:
180-x = 2(90+x)-36
180-x = 180+2x-36
180-x = 144+2x
3x = 180-144
3x = 36
x = 12 degrees (the angle)
180-x = 168 degrees (the supplement of the angle)
====================================================
cheers,
Stan H.